Status of all VLBA4 runs at AO since April 2001

Click on project # for the Tsys/gain info, and results / feedback etc.
Date
Project #
Network
P.I.
RX Used
Correlator
Brief Comment
26-Apr-01
TA025
VLBA+Ar
J. Romney
L-Narrow
Socorro
Fringes Found, X-pol, missing trk 14 and ~7--micron head-positioning error
15-Aug-01
TA026
VLBA+GBT+Ar
J. Romney
S-radar
Socorro
Fringes found after many trials which finally a 700 mu head positioner shift, all tracks Fine
07-Sep-01
TA027
VLBA+Ar
J. Romney
L-Narrow
Socorro
Fringes Found, with 700 micron shift. All tracks O.K.
15-Sep-01
BM159b
VLBA+Y+Ar
E. Momjian
L-Narrow
Socorro
Fringes Found after 700 micron + ~460 micron in the first pass shift. All tracks OK.
05-Oct-01
SC01C3
EVN+Ar
Tony Foley
C-band
JIVE
No Playback. This was a dual-head recording test. Some EVN stations had set-up problems. The run was rescheduled. AO couldn't join in.
New Headcal done with corrected procedure. 700 micron shift removed
18-Oct-01
W323D
VSOP+Ar+etc.
D.L. Jones
C-band
Socorro
Track#18 dead, ch2 dead for even passes. Fringes detected on Ch 3 and 4 (+ on ch2 in odd passes). Headstack enableing commands had errors which needed correcting online.
19-Oct-01
W323C
VSOP+Ar+etc.
D.L. Jones
C-band
Socorro
Same as above
09-Nov-01
ED018A
EVN+Ar
J.-F. DESMURS
Cb (5cm)
JIVE
Playback problems. A few EVN stations had freq-setting problems. No more news
10-Nov-01
ED018B
EVN+Ar
J.-F. DESMURS
Cb (5cm)
JIVE
--- no news ---
17-Nov-01
GP030A
Global
R. Porcas
C-band
Socorro
Good Fringes, all tracks/passes OK
21-Nov-01
EP034A
EVN+Ar
Y. Pihlstroem
L-wide
JIVE
Good Fringes, all tracks/passes OK
22-Nov-01
EP034B
EVN+Ar
Y. Pihlstroem
L-wide
JIVE
-- no news ----
04-Jan-02
BF063
VLA-Y,VLBA,Ar
J. D. Fix
L-wide
Socorro
Good Fringes, all tracks/passes OK
27-Jan-02
BC113B
VLBA+Ar
S. Chatterjee
L-wide
Socorro
No fringes (1pps signal jumpy, bad contact). Obs. is rescheduled for 19th April,02
04-Feb-02
F02l1
EVN+Ar
Test run
L-wide
JIVE
No fringes (1pps signal jumpy, bad contact)
20-Feb-02
TF022
VLBA+Ar
Test run
L-wide
Socorro
Fringes found, but no data read from track#18
24-Feb-02
GL026
GBT+Y+VLBA+Ar
C. Lonsdale
L-wide
Socorro
Tape accidentally erased before correlation. Will be rescheduled.
25-Feb-02
GP030B
Global
R. Porcas
L-wide
Socorro
Tape accidentally erased before correlation
19-Apr-02
BC113C
VLBA+Ar
S. Chatterjee
L-wide
Socorro
Fringes Found, both ungated and gated. Data recording was done via headstack#2 as track#18 on headstack#1 was unreliable.
12-May-02
BL106
VLBA+Y+GB+Ar
J. Lazio
L-wide
Socorro
Fringes Found, recording was headstack#2
30-May-02
N02C1
EVN+Ar
Z. Paragi
C-band
JIVE
Dual-head 512 Mbps, Network Monitoring Experiment. Recording was fine, but tracking troubles during the sort, 20 min run.
7-Jun-02
N02L2
EVN+Ar
Z. Paragi
L-wide
JIVE
Dual head, 512 Mbps recording, Network Monitoring Experiment at Lband. Data are being processed.
Track#18 problem fixed by Jon Hagen
19-Sep-02
TADU2
WSRT+Ar
Tony Foley
C-band
JIVE
Dual head, 512 Mbps recording, Testing WSRT
16-Nov-02
GL026B
Ar,EVN,VLBA,VLA-Y,GBT
Colin Lonsdale
L-wide
Socorro
High sensitivity mapping
17-Nov-02
N02L3
EVN+Ar
Z. Paragi
L-wide
JIVE
Network Monitoring Expt.
18-Nov-02
GM047
Ar,EVN,VLBA,GBT,Rb
J.M.Marcaide
L-wide
Socorro
Mapping
22-Nov-02
BS116A
VLBA-SC + Ar
S. Slysh
C-band
Socorro
Arecibo Holography
04-Feb-03
BM176
VLBA,Ar,VLA-Y
E. Momjian
L-narrow
Socorro
HI-absoprtion mapping
11-Mar-03
BC134A
VLBA+Ar
S Chatterjee
L-Wide
Socorro
VLB Proper Motion of a Faint Anomalously-Located Pulsar - Two channels correlated (freq-setting error).
02-Apr-03
BM176B
VLBA+Y27+Ar
E. Momjian
L-wide
Socorro
Continuum and 21cm HI Absorption toward Arp193 - correlated
18-Apr-03
R03N2
Ar
T. Ghosh
L-band
JIVE
2-head recording test (at 512 Mbps) Good quality data on all tracks
06-Jun-03
ED018C
EVN+Ar
J.-F. DESMURS
Cb (5cm)
JIVE
6GHz OH maser observations in Massives Star-Forming Regions - successful
07-Jun-03
ED018D
EVN+Ar
J.-F. DESMURS
Cb (5cm)
JIVE
6GHz OH maser observations in Massives Star-Forming Regions - successful
11-Jun-03
BS116B
Ar+VLBA-SC+HN
Slava Slysh
Cb
Socorro
Arecibo Holography - successful
20-Jun-03
BT069E
VLBA+EB+Y27+
GBT+Ar+WSRT
Greg Taylor
8.4 GHz
Socorro
4cm-mapping of GRB 030329 - this was the first X-band VLBI at Ar. Polarization handedness was reversed. Good correlated data after channel reversal.
17-Jul-03
BC134B
VLBA+Ar
S. Chatterjee
L-wide
Socorro
VLB Proper Motion of a Faint Anomalously-Located Pulsar -correlation under progress
31-Aug-03
BM193A
VLBA+GBT+Y27+Ar
E. Momjian
X-band
Socorro
LO was out of lock at Ar - no fringes
1-Sep-03
BM193B
VLBA+GBT+Y27+Ar
E. Momjian
C-band
Socorro
Good
11-Oct-03
V1609A
Ar+Eb+Wb+Jb
R. Porcas
L-band
Bonn
Good
1-Nov-03
GT005
EVN+VLBA+GBT+Ar+Y
G. Taylor
X-band
JIVE
Good
7-Nov-03
N03L3
EVN+Ar
Z. Paragi
L-band
JIVE
Good
9-Nov-03
GD017A
EVN+VLBA+GBT+Y+Ar
P. Diamond
L-band
Socorro
Good
2-Dec-03
BC134C
VLBA+Ar
S. Chatterjee
L-band
Socorro
Good
20-Mar-04
BB178A
VLBA+Ar+GBT
G. Bower
X-band
Socorro

20-Mar-04 V1609B Ar+Eb
R. Porcas
L-band Bonn Good
21-Mar-04 BB178B VLBA+Ar+GBT G. Bower X-band Socorro
21-Mar-04 V1609C Ar+Eb
R. Porcas L-band Bonn Good
21-May-04 TA031
VLBA+Ar
C. Salter
X, L-bands Socorro

3-Jun-04
N04L2
EVN+Ar
Z. Paragi
L-band JIVE

3-Jun-04
EG030
EVN+Ar
T. Ghosh
L-band
JIVE

9-Jun-04
GV017
EVN+VLBA+GBT+Ar
R. Vermeulen
L-band Socorro

24-Jun-04
V1609D Ar+Eb
R. Porcas L-band Bonn Good
26-Jun-04
V1609E Ar+Eb
R. Porcas L-band Bonn Good
21-Oct-04
GR024A
VLBA+EVN+GBT+Ar
R. Ramachandran
P-band
Socorro
Good
21-Oct-04
GR024B
VLBA+EVN+GBT+Ar
R. Ramachandran
P-band
Socorro

22-Oct-04
GG056
VLBA+EVN+Ar
C. Gwinn
P-band
Socorro

23-Oct-04
GG058B
VLBA+EVN+Ar C. Gwinn P-band Socorro
24-Oct-04
GG058C
VLBA+EVN+Ar C. Gwinn P-band Socorro
25-Oct-04
GR024C
VLBA+EVN+GBT+Ar
R. Ramachandran
P-band
Socorro

22-Nov-04 BB190
HSA
M. Bietenholz X-band
Socorro
Good
23-Dec-04
BH128
HSA
D. Hough
X-band
Socorro
Good
26-Dec-04
BB191
HSA
R. Barvainis C-band
Socorro
Good
27-Dec-04
BJ054
HSA
N. Jackson
L-band
Socorro
Good
25-Feb-05
GB053
VLBA+EVN+Ar
N. Bartel
C-band Socorro
26-Feb-05
BU030A
HSA
J. Ulvestad
C-band Socorro
7-Mar-05
GD017B
VLBA+EVN+GBT+Y27+Ar
P. Diamond
L-band
Socorro

29-May-05 BB204
HSA
A. Biggs
L-band
Socorro
10-Jun-05
BB209A
HSA E. Boyce
X-band Socorro
10-Jun-05 BW084
HSA J. Winn C-band Socorro
13-Jun-05 BM223A
HSA T. Maccaroni
C-band Socorro
3-Jul-05
BF084A HSA J. Forbrich
X-band Socorro Wrong freq setup at Ar
22-Jul-05
BB191B
HSA R. Barvainis C-band Socorro Cable Problem at Ar
24-Jul-05
BM223B
HSA T. Maccaroni C-band Socorro
28-Jul-05
BF084B
HSA J. Forbrich X-band Socorro
19-Aug-05
BB209B
HSA
E. Boyce C-band Socorro
21-Aug-05
BJ059
HSA
P. Jonker C-band Socorro
4-Sep-05
BB203B
HSA
R. Barvainis C-band Socorro
4-Sep-05 BB203C HSA R. Barvainis C-band Socorro
9-Dec-05
BB217a
HSA
E. Boyce C-band Socorro
30-Jan-06
BB191C
HSA J. Ulvestad
C-band Socorro
2-Mar-06
GM062A
VLBA+EVN+GBT+Ar M. Orienti
L-band
JIVE

26-Mar-06 BH139
HSA
D. Hough X-band Socorro VLBA4 recorder track separation problem

For a list of Mark 5 runs, click here.


Last updated: Friday, 28-Apr-2006 16:00:40 AST
E. Momjian